A particle is displayed 60 cm to the north and then is given a second displacement of 90 cm to the east. what is the average velocity of the particle if it covers the displacement within 2.15 


1
Expert's answer
2021-10-14T11:36:37-0400

First, find the displacement:


"d=\sqrt{0.6^2+0.9^2}=1.08\text{ m}."

The average velocity is the displacement over time:


"v=\frac{d}{t}=\frac{1.08}{2.15}=0.5\text{ m/s}."